Prove that in a graph G, the sum of the degrees of the vertices is equal to twice the number of edges. Consequently, the number of vertices with odd degree is even

1
Expert's answer
2021-11-24T15:52:09-0500

Let G be a graph with e edges and n vertices v1,v2,v3,...,vn.

Since each edge is incident on two vertices, it contributes 2 to the sum of degree of vertices in graph G. Thus the sum of degrees of all vertices in G is twice the number of edges in G:"\u2211^n_{i=1}degree(v_i)=2e"


Let the degrees of first r vertices be even and the remaining (n−r) vertices have odd degrees, then:

"\u2211^n_{i=1}degree(v_i)=\u2211^r_{i=1}degree(v_i)+\u2211^n_{i=r+1}degree(v_i)"


"\implies ∑^n_{i=1}degree(v_i)+∑^r_{i=1}degree(v_i)"


"\implies ∑^n_{i=r+1}degree(v_i)" is even.


But, the for "i=r+1,r+2,...,n" each "d(v_i)" is odd. So, the number of terms in

"\u2211^n_{i=r+1}degree(v_i)" must be even. So, "(n-r)" is even.